OpenStudy (anonymous):

parametric question: Find the PQ for y if x = e^t for the line passing through the points (2,1) and (-2,3). I know basic parametrics, but how do I solve a question like this??

OpenStudy (dumbcow):

x = e^t ----> t = ln(x) the point (-2,3) cant exist since "x" is negative ??

OpenStudy (loser66):

the line passes through the points (2,1) and (-2,3) has equation \(y= 2-\dfrac{1}{2}x\), replace \(x=e^t\) you have the parametric equation for y is \(y= 2-\dfrac{e^t}{2}\) DAT SIT

OpenStudy (anonymous):

Ah... Thanks a lot!